### v1

# Optimization of Module Transferability in Single Image Super-Resolution: Universality Assessment and Cycle Residual Blocks

Abstract
Deep learning has substantially advanced the Single Image Super-Resolution (SISR). However, existing researches have predominantly focused on raw performance gains, with little attention paid to quantifying the transferability of architectural components. In this paper, we introduce the concept of “Universality” and its associated definitions which extend the traditional notion of “Generalization” to encompass the modules’ ease of transferability, thus revealing the relationships between module universality and model generalizability. Then we propose the Universality Assessment Equation (UAE), a metric for quantifying how readily a given module could be transplanted across models. Guided by the UAE results of standard residual blocks and other plug-and-play modules, we further design two optimized modules, Cycle Residual Block (CRB) and Depth-Wise Cycle Residual Block (DCRB). Through comprehensive experiments on natural-scene benchmarks, remote-sensing datasets, extreme-industrial imagery and on-device deployments, we demonstrate that networks embedded with the proposed plug-and-play modules outperform several state-of-the-arts, reaching a PSNR enhancement of up to 0.83dB or enabling a 71.3% reduction in parameters with negligible loss in reconstruction fidelity.
Keywords Generalization; Optimization; Super-resolution; Universality
1 Introduction
Single Image Super-Resolution (SISR) reconstructs high-resolution images from low-resolution inputs. However, as application scenarios expand, deep SISR methods increasingly exhibit limited generalization. Therefore, numerous intricately designed architectures have been proposed to enhance model performance across diverse scenarios.
Since Dong et al.[1] firstly introduced deep learning methods into image super-resolution tasks, deep learning-based SISR methods have gradually become mainstream. Methods such as SRCNN [1], FSRCNN (Fast SRCNN) [2], and ESPCN (Efficient Sub-Pixel CNN) [3] harness the learning capabilities of shallow convolutional networks, leading to notable performance improvements over traditional methods. Subsequently, models such as VDSR (Very Deep CNN for SR) [4], RED-Net (Residual Encoder-Decoder Network) [5], and EDSR (Enhanced Deep SR) [6] began exploring deeper architectures to improve the SISR performance, among which Residual Blocks (RB) have become a commonly used component. For instance, EDSR cascades several residual modules that remove Batch Normalization (BN) layers to serve as the feature extraction layer. MSRN (Multi-Scale Residual Network) [7] introduces convolution kernels of different sizes within residual modules. Residual blocks and their optimized variants have become one of the fundamental components in SISR networks.
Inspired by DenseNet, several SISR networks based on dense convolution modules are also proposed. SRDenseNet [8] builds the feature inference blocks by cascading dense connection modules and skip connections between different layers. RDN (Residual Dense Network) [9] combines residual modules and dense connection modules to form a residual-dense network. MemNet [10] constructs a memory module through dense connections between different convolution layers and gating units to maintain long-term dependencies between features at different layers.
In recent years, various attention modules have been proposed to improve the model performance. For instance, RCAN [11] introduces channel attention into residual modules and builds a dual-layer residual structure to enhance feature inference capabilities. SAN [12] proposes second-order channel attention and constructs non-local enhanced residual groups. Additionally, Transformer-based architectures have also demonstrated considerable potential in SISR tasks. SwinIR (Image Restoration Using Swin Transformer) [13] achieves better results with fewer parameters by constructing a Residual Swin Transformer module. HiT-SR [14] builds a Hierarchical Transformer module that expands windows to aggregate hierarchical features, enabling feature inference from local regions to long-range dependencies.
It is evident that various modules form the core components of SISR models, including residual modules, Transformer modules, etc. Additionally, researchers increasingly favor the plug-and-play modules characterized by simplicity in structure, ideal performance and strong transferability. For instance, ConvFFN was initially introduced in ViT [16] and is still applied in IPG [17]. ResBlock, primarily proposed in [18], is still widely embedded in various networks to enhance model performance. In existing studies, however, the plug-and-play modules are characterized only qualitatively, and the concept of “Generalization” primarily focuses on performance across different datasets. Although numerous state-of-the-arts have markedly improved the SISR performance, their internal modules remain difficult to operate independently. For instance, although IPG [17] achieves ideal reconstruction performance using a delicately optimized GNN architecture, its modules are highly coupled–a limitation also observed in SwinIR [13], as shown in Fig.1.
Figure 1: The highly coupled internal structure of IPG (CVPR 2024) and SwinIR (ICCV 2023).
To enable a quantitative assessment of module transferability and provide a novel perspective on plug-and-play module optimization, this paper proposes the definition of “Universality” and its associated concepts. Subsequently, the Universality Assessment Equation is designed by analyzing the pivotal characteristics of existing plug-and-play modules and other state-of-the-art networks. Cycle Residual Blocks (CRB) and Depth-wise Cycle Resiual Blocks (DCRB) are further derived on the basis of RB under the guidance of UAE results. Experiments demonstrate that models embedded with the optimized modules could achieve a 0.83dB increase in PSNR, or reduce parameters by up to 71.3% with minimal performance loss. The main contributions of this paper are as follows:
1) The definition of “Universality” is proposed. Compared to “Generalization”, “Universality” describes modules’ ease of transferability and is quantified by the Universality Assessment Equation (UAE).
2) CRB and DCRB are designed as optimized plug-and-play modules referring to UAE results. The inherent optimization mechanisms of CRB and DCRB are also revealed from the standpoint of back propagation process.
3) The effectiveness of CRB and DCRB is validated through experiments on various SISR tasks ranging from conventional settings to extreme industrial scenarios (e.g., coalmine images), where the optimized modules outperform several state-of-the-arts and significantly enhance the visual quality of reconstructed images. Moreover, through on-device deployments, we have validated the practicability of CRB in real-world applications.
2 Related Works
2.1 Models for Single Image Super-Resolution
Existing SISR networks could be broadly classified into four categories based on their backbone models: CNN-based models ([1, 6, 4]), GNN-based models ([17, 20, 21, 22]), GAN-based models ([23]), and Transformer-based models ([15, 13, 42, 24]). While highly integrated models could reduce computational costs and improve the performance, they also make it difficult to achieve efficient plug-and-play compatibility across different networks. For instance, the RSTB module in SwinIR [13] contains STL layers, which in turn incorporate MSA (Multi-Head Self-Attention) and MLP (Multi-Layer Perceptron). The deep integration of these components requires careful structure modifications when transferring the module to other backbones.
2.2 Plug-and-Play Modules in SISR Networks
Existing SISR methods often incorporate various modules into the networks to enhance model performance. For instance, RB, introduced in ResNet [18], has been widely adopted in diverse architectures. ConvFFN, introduced in ViT [16], is used in the GNN-based IPG [17]. SRFlow-DA [27] embeds RB directly into the affine injector layer to improve the performance of SRFlow [28]. These plug-and-play modules significantly enhance model performance but are relatively scarce and difficult to design, since structures of the plug-and-play modules are required to be highly streamlined while preserving an ideal performance. Therefore, we attempt to analyze the structure of plug-and-play modules from a universality perspective, thus designing modules with enhanced performance.
2.3 Generalization and Universality
“Generalization” denotes a model’s ability to maintain high performance on test sets distinct from its training data [29]. IPT [15], for example, utilizes large-scale pre-training to unify three different low-level vision tasks. DGUNet [19], based on DNN, designs an interpretable SISR network featuring universality. While these studies enhance model performance by improving generalizability, they lack a quantitive evaluation of local module transferability. Therefore, this paper introduces the concept of universality to characterize module transferability, thereby facilitating the design of plug-and-play modules that enhance model generalization.
3 Proposed Methods
In this section, definitions for “Universality” and its associated concepts are proposed to lay the foundation for a qualitative description of module transferability, namely the Universality Assessment Equation (UAE). We further optimize RB into CRB/DCRB through UAE analysis and the optimization mechanism under UAE guidance is elucidated from the perspective of back propagation. Finally, a generalized designing principle for CRB is explained to provide a clearer picture on module optimization.
3.1 “Universality” and “Positive Universality”
Let $M_{i}$ and $M_{k}$ denote two network architectures ($i\neq k$), and let $B_{j}$ be a module originally embedded in $M_{i}$. We define $B_{j}$ to be universal for $M_{k}$ if it could be incorporated into $M_{k}$ without significant structural modifications while preserving its intended functionality. Thus, universality measures modules’ ease of transferability across architectures. By contrast, generalization refers to a model’s ability to maintain an ideal performance when evaluated on datasets distinct from its training set. A summary of these distinctions is presented in Table 1.
Table 1: Differences between “Universality” and “Generalization”.
Concept Subject Object Evaluation Criteria
Universality Local module Complete model The ease of module transferability
Generalization Complete model Brand new test data Whether model performance remains ideal
Although universality could describe a module’s intrinsic transferability, it is difficult to characterize its precise impacts on model performance after integration, where directly embedding a plug-and-play module may degrade the performance. For instance, inserting a PnP denoising block [30] into the EDSR backbone is likely to trigger performance degradation, because the denoiser’s learned feature distribution breaks the identity mappings that deep residual blocks rely on, which perturbs the signals. Therefore, we propose the “Positive Universality (PU)” to describe that module $B_{j}$ could effectively improve the performance of a brand new model after being detached from its parent model. The opposite concept is “Negative Universality (NU)”. Our experiments validate that the CRB is such a module possessing PU property.
Universality and generalization are not completely isolated. Differences between the previously proposed plug-and-play modules ([18, 11, 31, 32]) reveal that while more complex blocks could achieve superior generalizability, they also typically incur larger parameter counts and more elaborate computational graphs, all of which may deteriorate the ease of transferability. In other words, model’s generalizability could be enhanced by appropriately decreasing the module universality.
3.2 Universality Assessment Equation (UAE)
Through the structural differences between plug-and-play modules (RB [18], RCAB [11], ConvFFN [16], etc.) and non-plug-and-play modules (RSTB [13], GAL [17], etc.), we identify four shared factors that influence most on the universality, namely “The nesting level of other blocks”, “Total number of parameters”, “Cascading of forward propagation layers” and “Input feature dimensions”. Six UAE forms are chosen in Experimental Analysis to validate the conclusions.
Generally, transferability difficulty rises with an increased module nesting level due to stronger internal coupling (e.g., GAL [17]). Parameter-scarce modules are sensitive to parameter growth until a threshold, beyond which further increases yield diminishing returns and a risk of overfitting (e.g., standard convolution modules [1] and RB [18]). Augmenting the number of forward propagation layers increases computational overhead and correspondingly deteriorates module universality. However, this degradation is less pronounced than the increases contributed by nested architectures, since the nested modules already incorporate multiple propagation stages (e.g., ConvFFN [16] and RSTB [13]). Consequently, we compute the UAE per input feature-map unit, defining it abstractly as:
$\phi=\frac{\alpha(l)\times\beta(k)\times\theta(n)}{f},\ \begin{cases}\theta^{\prime\prime}(n)<0,\\ \beta^{\prime\prime}(k)>\alpha^{\prime\prime}(l)>0.\end{cases}$ (1)
where $k$ is the number of nested sub-modules, $n$ the total parameters, $l$ the forward propagation layers, $f$ the input feature units. $\phi$ is inversely related to universality. $\theta^{\prime\prime}(n)<0$ denotes a diminishing sensitivity to parameter increases and a slowing universality decline, while $\beta^{\prime\prime}(k)>\alpha^{\prime\prime}(l)>0$ indicates that nesting modules erode universality more rapidly than cascaded layers.
Although the specific forms of $\alpha(l),\beta(k)$ and $\theta(n)$ vary, the choice of UAE does not alter the relative ranking of module universality. Eq.(2) provides one such instance, and five more UAE variants —despite spanning from $10^{-1}$ to $10^{4}$—yield identical module orderings. Consequently, modules with UAE values close to the baseline exhibit similar transferability, and lower UAE results generally indicate better universality.
$\phi=\frac{l\times e^{k+1}\times\lg\left(\frac{n}{100}\right)}{f}$ (2)
Figure 2: Structures of the proposed CRB and DCRB, where $l$ stands for the UAE variable.
3.3 Cycle Residual Block and Depth-wise Separable Cycle Residual Block
According to Eq.(1), RB could be optimized from three perspectives: $k$, $l$, and $n$. This study optimizes RB with respect to $l$ and $n$, yielding CRB and DCRB respectively, as illustrated in Fig.2. Specifically, CRB reuses RB’s standard convolution layers, thereby improving the layer-utilization efficiency without adding any extra parameter. DCRB employs depth-wise separable convolution to reduce the parameters, resulting in a more lightweight and integrated module structure.
Let $X\in{R}^{B\times C\times H\times W}$ denote the input tensor, where $B$, $C$, $H$ and $W$ correspond to batch size, channel dimensions, height and width, respectively. The CRB and its depth-wise variant DCRB are then defined as Eq.(3), where Conv($\cdot$) represents standard convolution for CRB and depth-wise separable convolution for DCRB.
$\begin{cases}X_{\text{mid}}=X_{\text{input}}+\text{Conv}(\text{ReLU}(\text{Conv}(X_{\text{input}})))\\ Y_{\text{output}}=X_{\text{mid}}+\text{Conv}(\text{ReLU}(\text{Conv}(X_{\text{mid}})))\end{cases}$ (3)
To elucidate the optimization mechanism of CRB, we further examine its back propagation process. Initially, to assume that gradient magnitude could be taken as a approximation for convergence rate, we enforce several prior conditions of RB and CRB to be identical, all of which are ensured in our experiments via the following configurations: (1) We use identical hyperparameters (learning rate, etc.), (2) The BatchNorm layers of RB/CRB are removed, and the global random seeds are fixed to control the identical condition number $\kappa=L/\mu$, (3) We apply the same dropout rate and other settings to ensure similar local Hessian structures. Under these controlled conditions, the observed difference in gradient magnitude reflects the difference of per-step update, which could be interpreted as the convergence speed.
Let $y=f_{L}(f_{L-1}(\cdots f_{l}(x)\cdots))$ denote a neural network where $x$ is the input of layer $f_{l}(\cdot)$ and $y$ the model output, the back propagation is then given by Eq.(4).
$\frac{\partial L(x)}{\partial x}=\frac{\partial L(x)}{\partial y}\cdot\frac{\partial y}{\partial x}=\frac{\partial L(x)}{\partial y}\cdot\prod_{i=l}^{L-1}\frac{\partial x_{i+1}}{\partial x_{i}}=\frac{\partial L}{\partial y}\cdot\prod_{i=l}^{L-1}J_{i}$ (4)
where $L(x)$ stands for losses and $J_{i}$ the Jacobian matrix. Let $\mathbf{F}_{l}(\cdot):\mathbb{R}^{d}\xrightarrow{}\mathbb{R}^{d}$ denote the nonlinear mapping (e.g., convolution followed by activation ) in the $l$-th residual block. Since each block implements $x_{l+1}=x_{l}+\mathbf{F}_{l}(x_{l})$, the total derivative of the loss with respect to initial input $x_{0}$ follows by the repeated application of chain rule:
$\frac{\partial L(x)}{\partial x_{0}}=\frac{\partial L(x)}{\partial x_{n}}\cdot\prod_{l=0}^{n-1}(\mathbf{E}+\frac{\partial\mathbf{F}_{l}(x_{l})}{\partial x_{l}})=\frac{\partial L(x)}{\partial x_{n}}\prod_{l=0}^{n-1}J_{l}^{\text{RB}}$ (5)
where $\textbf{E}\in\mathbb{R}^{d\times d}$ is the identity matrix. From Eq.(3), CRB could be expressed as :
$y=x+\mathbf{F}(x)+\mathbf{F}(x+\mathbf{F}(x))$ (6)
By conducting similar calculations, we derive the back propagation of a model made up of $n$ CRBs:
$\frac{\partial L(x)}{\partial x_{0}}=\frac{\partial L(x)}{\partial x_{n}}\prod_{l=0}^{n-1}\Big{[}(\mathbf{E}+\frac{\partial\mathbf{F}_{l}(x_{l})}{\partial x_{l}})(\mathbf{E}+\frac{\partial\mathbf{F}_{l}(y_{l})}{\partial y_{l}})\Big{]}=\frac{\partial L(x)}{\partial x_{n}}\prod_{l=0}^{n-1}J_{l}^{\text{CRB}}.$ (7)
where $y_{l}=x_{l}+\mathbf{F}_{l}(x_{l})$. Considering that $\mathbf{E}+\partial\mathbf{F}_{l}(x_{l})/\partial x_{l}$ are matrices, we explicitly employ the spectral norm $\|\cdot\|_{2}$ to characterize the gradient amplification. Firstly, it is proved that when $A$ and $B$ are symmetric positive semidefinite matrices and commute with the identity matrix, Eq.(8) is valid.
$\|(E+A)(E+B)\|_{2}>\|(E+A)\|_{2}.$ (8)
Figure 3: Experimental validation of Eqs.(5), (7) and (12), where EDSR [6] is chosen as the backbone. EDSR$+$CRB attains 31.00 dB PSNR by epoch 200, whereas EDSR$+$RB achieves the same value around epoch 350, confirming CRB’s faster convergence.
Figure 4: Experimental validation of Eqs.(14) and (16), where EDSR [6] is chosen as the backbone. CRBl=16,ε=4 exhibits smooth convergence, whereas CRBl=16,ε=5 fails to converge and possesses a substantially degraded performance, with its loss plateauing at approximately 632.5 during the experiments.
Let $A,B\in\mathbb{R}^{d\times d}$ satisfy: (1) “$A$ and $B$ are symmetric positive semidefinite matrices”, (2) “$A$ and $B$ commute, i.e. $AB=BA$”. Thus, there exists orthogonal matrix $Q$ satisfying Eq.(9).
$Q^{\top}AQ=\text{diag}(\alpha_{1},\cdots,\alpha_{d}),Q^{\top}BQ=\text{diag}(\beta_{1},\cdots,\beta_{d}).$ (9)
where $\alpha_{i},\ \beta_{i}>0$. The matrices $E+A$ and $E+B$ are diagonalized under this basis as Eq.(10).
$Q^{\top}(E+A)Q=\text{diag}(1+\alpha_{1},\cdots,1+\alpha_{d}),Q^{\top}(E+B)Q=\text{diag}(1+\beta_{1},\cdots,1+\beta_{d}).$ (10)
Thus, we have $(E+A)(E+B)$ in this basis equal to $\text{diag}((1+\alpha_{1})(1+\beta_{1}),\cdots,(1+\alpha_{d})(1+\beta_{d}))$. By the definition of $\|\cdot\|_{2}$, we derive Eq.(11).
[…]
$\|(E+A)(E+B)\|_{2}$ $=(1+\alpha_{\max})(1+\beta_{\max})>(1+\alpha_{\max})$ $=\|(E+A)\|_{2}$ (12)
Although the proof relies on strong linear algebra assumptions where $J_{l}^{\text{RB}}$, $J_{l}^{\text{CRB}}$ may not be symmetric, and its eigenvectors may not be identical, the core insight still holds: the residual structure inherently inserts a gain factor $\geq 1$ into the chain product, thereby revealing how residual structures mitigate gradient vanishing/explosion.
Such approximation is grounded in existing physical principles. From the perspective of Neural-ODEs [33], a residual block implements one step of the Euler method: $x(t+\Delta t)=x(t)+\Delta t\cdot f(x(t))$, yielding $F(x)=x(t+\Delta t)-x(t)\approx\Delta t\cdot f(x)$ and $F^{\prime}\xrightarrow{}0$. Treating $F^{\prime}$ as a “small perturbation” aligns with the fact that residual learning ultimately guides deep modules degenerate into identity function. In quantum mechanics, Dysen series and Lie-Trotter product formulas [34] use the approximation $e^{\frac{A}{n}}\approx I+\frac{1}{n}A+O(n^{-2})$ retaining merely the linear “$I+$small” at each step, motivating the requirements $\|E+A\|_{2}\geq 1$ in Eq.(12).
Therefore, even relaxing the symmetric/commuting conditions, Eq.(12) together with Eq.(7) indicate that every added residual branch contributes at least a unit of spectral gain1absent1$\geq$ 1≥ 1 into the back propagation, leading to smoother gradient flow, faster convergence, and better module performance, as is verified through experiments in Fig.3.
Table 2: UAE calculation on different modules, where $\text{UAE}_{i}$ represents six specific forms of UAE and $f=64$. Structure of CRB is chosen as $\varepsilon=2,l=8$.
RB (baseline) [18] RCAB [11] ConvFFN [16] RSTB [13] GAL [17] DCRB CRB
UAE Variables $k$ 0 1 1 3 3 1 0
$n$ 73,856 148,292 17,856 86,784 56,132 1,280 73,856
$l$ 4 15 6 11 21 8 8
$\text{UAE}_{1}(\times 1)$ $\phi_{1}$ 0.49 5.50 1.56 27.57 49.25 1.02 0.98
$\text{UAE}_{2}(\times 1)$ $\phi_{2}$ 0.17 1.73 0.69 9.38 17.92 0.92 0.34
$\text{UAE}_{3}(\times 1)$ $\phi_{3}$ 0.13 1.31 0.37 1.98 3.54 0.24 0.26
$\text{UAE}_{4}(\times 1)$ $\phi_{4}$ 1.95 82.38 9.36 303.32 1034.29 8.18 7.80
$\text{UAE}_{5}(\times 1)$ $\phi_{5}$ 72.19 8159.08 463.54 25832.26 75268.47 278.18 276.01
$\text{UAE}_{6}(\times 1)$ $\phi_{6}$ 0.08 0.71 0.29 1.20 2.30 0.38 0.16
Table 3: Quantitative comparison results (PSNR/SSIM) between several universal modules with 4×4$\times$4 × SR scale, trained on DIV2K. Bold texts indicate the best performance.
Modules RB (baseline) [18] RCAB [11] ConvFFN [16] CRB DCRB
$\phi_{2}$ 0.17 1.73 0.69 0.34 0.92
Set5 [40] 31.29/0.8632 31.17/0.8542 31.71/0.8954 31.33/0.8648 30.91/0.8538
Set14 [39] 28.06/0.6961 27.07/0.7089 27.86/0.7088 28.49/0.6971 26.69/0.6907
B100 [36] 25.95/0.8122 25.89/0.8082 25.91/0.8109 25.97/0.8137 25.89/0.8076
Urban100 [41] 23.53/0.7597 23.49/0.7414 23.45/0.7502 23.52/0.7615 23.13/0.7312
[…]
SR Scale Model Set5[40] Set14 [39] B100 [36] Urban100 [41]
×$\times 3$ EDSR (RB) 33.69/0.8844/0.0707 29.87/0.7101/0.0634 28.31/0.8540/0.0980 -/-/-
Ours1(CRB) 33.76/0.8846/0.0684 29.92/0.7114/0.0624 28.33/0.8553/0.0972 -/-/-
Ours2(DCRB) 32.98/0.8756/0.1248 28.38/0.7072/0.0918 27.56/0.8421/0.1538 -/-/-
×$\times 4$ EDSR (RB) 31.29/0.8632/0.1092 28.06/0.6961/0.1085 25.95/0.8122/0.2026 23.53/0.7597/0.0247
Ours1(CRB) 31.32/0.8648/0.1112 28.49/0.6971/0.1110 25.97/0.8137/0.2008 23.52/0.7615/0.0243
Ours2(DCRB) 30.91/0.8538/0.2017 26.69/0.6907/0.1690 25.89/0.8076/0.2853 23.13/0.7312/0.0236
3.4 Generalized Forms for CRBs and its Designing Principles
Although Eq.(12) demonstrates that introducing extra residual connections accelerates module convergence without adding extra parameters, we derive an upper bound of $\frac{l}{4}$ on the residual number $\varepsilon$ ($\varepsilon\leqslant\frac{l}{4}$), where $l$ is the CRB’s UAE variable and the constant 4 equals half the number of forward propagation layers per CRB.
Following Eq.(7), we initially verify that for $\varepsilon=4,l=16$, CRB converges stably. From Fig.2, CRB ($\varepsilon=4,l=16$) could be expressed as Eq.(13)
$y_{1}=x+\mathbf{F}_{1}(x),y_{2}=y_{1}+\mathbf{F}_{2}(y_{1}),y_{3}=y_{2}+\mathbf{F}_{3}(y_{2}),y_{4}=y_{3}+\mathbf{F}_{4}(y_{3}).$ (13)
Hence, for a network comprised of $L$ CRBε=4, the back propagation is given by Eq.(14). Fig.LABEL:converge_1 and Fig.LABEL:loss_1 provides the experimental validation for the convergence of Eq.(14).
$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}\Big{[}(\mathbf{E}+\frac{\partial\mathbf{F}_{4,l}}{\partial y_{3,l}})(\mathbf{E}+\frac{\partial\mathbf{F}_{3,l}}{\partial y_{2,l}})(\mathbf{E}+\frac{\partial\mathbf{F}_{2,l}}{\partial y_{1,l}})(\mathbf{E}+\frac{\partial\mathbf{F}_{1,l}}{\partial x_{l}})\Big{]}=\frac{\partial L}{\partial x_{L}}\prod_{i=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon=4}}\end{aligned}.$ (14)
However, when an additional residual connection is introduced to Eq.(13), as shown in Eq.(15), the back propagation should be updated as Eq.(16).
$y_{1}=y_{0}+\mathbf{F}_{1}(y_{0}),y_{2}=y_{1}+\mathbf{F}_{2}(y_{1})+x,y_{3}=y_{2}+\mathbf{F}_{3}(y_{2}),y_{4}=y_{3}+\mathbf{F}_{4}(y_{3}).$ (15)
$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\Bigg{[}\prod_{l=0}^{L-1}\Big{[}\prod_{i=1}^{4}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})+\Theta_{l}\Big{]}\Bigg{]}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon=5}},\Theta_{l}=\prod_{t=3}^{4}(\mathbf{E}+\frac{\partial\mathbf{F}_{t,l}}{\partial y_{t-1,l}}).$ (16)
In conjunction with Eq.(12), it is clear that $\|J_{l}^{\text{CRB}_{\varepsilon=5}}\|_{2}>2>\|J_{l}^{\text{CRB}_{\varepsilon=4}}\|_{2}>1$, indicating that $\text{CRB}_{l=16,\varepsilon=5}$ is highly susceptible to trigger gradient exploding. Such conclusion derived from theoretical analysis are corroborated by the experimental results presented in Fig.LABEL:converge_2 and Fig.LABEL:loss_2, where $\text{CRB}_{l=16,\varepsilon=5}$ exhibits gradient explosions in first 20 epochs accompanied by a substantial decline in PSNR performance.
We now generalize the setting to a CRB with $l$ forward propagation layers where $\varepsilon$ residual connections are introduced, as shown in Eq.(17). Assuming that the previously stated hypothesis on the upper bound of residual numbers $\varepsilon\leqslant\frac{l}{4}(\varepsilon\in\mathbb{N})$ holds, then the computed results should theoretically indicate that the module converges for $\varepsilon=\frac{l}{4}$ but fails to converge for $\varepsilon=\frac{l}{4}+1$.
$\begin{cases}y_{1}=x+\mathbf{F}_{1},\cdots,\ y_{\varepsilon}=y_{\varepsilon-1}+\mathbf{F}_{\varepsilon},&\varepsilon=\frac{l}{4}\\ y_{1}=x+\mathbf{F}_{1},\ y_{2}=y_{1}+\mathbf{F}_{2}+x,\cdots&\varepsilon=\frac{l}{4}+1\end{cases}$ (17)
Referring to Eqs.(14) and (16), we derive the back propagation process ($\varepsilon=\frac{l}{4}$ for Eq.(18) and $\varepsilon=\frac{l}{4}+1$ for Eq.(19)) for the model made up of $L$ generalized CRBs. Similar to Eq.(12), it could be easily concluded that $\|J_{l}^{\text{CRB}_{\varepsilon=l/4+1}}\|_{2}>2>\|J_{l}^{\text{CRB}_{\varepsilon=l/4}}\|_{2}>1$. Consequently, when $\varepsilon>\frac{l}{4}$, $\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon}}$ exhibits exponential growth with a base larger than 2, leading to gradient explosion and declined module performance.
$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L}\prod_{i=1}^{\varepsilon}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon}}.$ (18)
$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\Bigg{[}\prod\limits_{l=0}^{L-1}\Big{[}\prod\limits_{i=1}^{\varepsilon}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})+\Theta_{l}\Big{]}\Bigg{]}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon}},\Theta_{l}=\prod_{t=3}^{\varepsilon}(\mathbf{E}+\frac{\partial\mathbf{F}_{t,l}}{\partial y_{t-1,l}}).$ (19)
Table 5: Quantitative results on Potsdam remote sensing dataset [43] under ×$\times 4$ SR scale. Hyper-parameter settings of IPG and SRFormer are scaled down to accommodate limited computational resources and accelerate training, while EDSR and SRResnet–both featuring moderate parameter counts–retain their full hyper-parameter configurations. The first 700 images of the dataset are used for model training, while the remaining 300 images for performance evaluation.
Models PSNR (dB) SSIM LPIPS
SRCNN (ECCV 2014)[1] 29.61 0.8572 0.2573
SRResnet (CVPR 2017) [18] 31.75 0.8986 0.0606
RCAN (ECCV 2018) [11] 31.98 0.9019 0.0593
EDSR (CVPR 2017) [6] 32.18 0.9042 0.0581
SRFormer (ICCV 2023)[42] 32.24 0.9054 0.0548
IPG (CVPR 2024)[17] 32.43 0.9058 0.0516
SRFormer (with CRB) 32.48 0.9062 0.0537
EDSR (with CRB) 32.49 0.9097 0.0526
EDSR (with DCRB) 32.33 0.8888 0.0543
SRFormer (with DCRB) 29.40 0.9041 0.0722
IPG (with CRB) 32.96 0.9049 0.0512
IPG (with DCRB) 32.85 0.9029 0.0565
Table 6: Quantitative results on BelT coalmine dataset [44] under ×$\times 4$ SR scale. Bold texts indicate the best performance.
Models PSNR(dB) SSIM LPIPS
EDSR (CVPR 2017) [6] 28.06 0.7551 0.1154
SRFormer (ICCV 2023) [42] 29.76 0.7906 0.1017
IPG (CVPR 2024) [17] 29.95 0.7921 0.1028
EDSR (with CRB) 28.54 0.7604 0.1122
IPG (with CRB) 30.03 0.7930 0.1010
IPG (with DCRB) 29.97 0.7928 0.0999
SRFormer (with CRB) 29.81 0.7884 0.1042
SRFormer (with DCRB) 29.80 0.7913 0.1033
In summary, the generalized CRB converges stably and achieves optimal performance only when $\varepsilon\leqslant\frac{l}{4}$ is satisfied, where $\varepsilon$ is the number of residual connections and $l$ the number of UAE propagation layers (Eq.(1)).
4 Experimental analysis
4.1 Experimental Setup and Implementation details
All experiments are conducted on an RTX 4090D GPU using the PyTorch framework. Training is performed on the DIV2K[35], BelT[44], Potsdam[43], with evaluation on B100, Set14, Set5, and Urban100 [36, 39, 40, 41]. Reconstruction fidelity is quantified by PSNR and SSIM [37], while perceptual quality is assessed via LPIPS [38], as shown in Eq.(20).
$\begin{cases}\text{PSNR}(x,y)=10\log_{10}(\frac{\text{MAX}^{2}}{\text{MSE}(x,y)})\\ MSE(x,y)=\frac{1}{MN}\sum_{i=0}^{M-1}\sum_{j=0}^{N-1}[x(i,j)-y(i,j)]^{2}\\ SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}\\ LPIPS(I_{1},I_{2})=\sum_{l}\lambda_{l}\|\phi_{l}(I_{1})-\phi_{l}(I_{2})\|_{2}\end{cases}$ (20)
where MAX is the maximum pixel value, MSE is the mean squared error between the ground truth and reconstructed image. $\mu_{1}$ and $\mu_{2}$ are the local mean intensities of the ground-truth and reconstructed images, $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$ represent their corresponding local variances. $\phi_{l}$ extracts the feature maps at the $l$-th layer of a pretrained network, and $\lambda_{l}$ denotes the learned weights. During the training of EDSR and the proposed variants, we adopt the $\mathcal{L}_{1}$ loss and ADAM optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$). The initial learning rate is set to $10^{-5}$.
When retraining baseline IPG [17] and SRFormer [42], we adjust hyperparameters to accelerate convergence. Specifically, the number of iterations is reduced to 10,000, the MLP ratio is lowered from 4 to 2, IPG’s embedding dimension is reduced to 20 with the number of heads set to 4 accordingly, and SRFormer’s dimension is set to 64. The core architectures of models, however, are still fully preserved, ensuring that the designing characteristics of networks remain unaffected despite the reductions in parameter size.
4.2 Universality Assessment for Diverse Modules
To validate Eq.(1) and demonstrate that varying the form of UAE does not alter module universality rankings, we evaluate six distinct UAEs (Eq.(21)) on RB [18], RCAB [11], ConvFFN [16], RSTB [13], GAL [17], CRB and DCRB. The resulting universality scores are reported in Table 2 and illustrated in Fig.7.
From Fig.7, although the absolute values differ markedly due to the varying nature of each equation, the relative ordering remains invariant: $\phi(\text{GAL})>\phi(\text{RSTB})>>\phi(\text{RCAB})>\phi(\text{ConvFFN})>\phi(\text{DCRB})\approx\phi(\text{CRB})>\phi(\text{RB})$ where RB is adopted as the reference baseline. Hence, CRB and DCRB attain universality comparable to RB. Moreover, the intermediate magnitudes and consistent ranking of $\phi_{1}$ and $\phi_{2}$ render analyses based on these two UAEs particularly transparent and interpretable.
Table 2 demonstrates the core determinants of universality. Although RCAB [11] possesses larger $n$ and $l$ compared to RSTB [13], the elevated $k$ in RSTB deteriorates its transferability. This finding aligns with empirical evidence: RSTB integrates three nested submodules—MSA, MLP, and STL—whereas RCAB employs only a CALayer. Hence, effective migration of RSTB requires the concurrent adaptation of multiple components. Conversely, despite $n_{\mathrm{CRB}}\gg n_{\mathrm{DCRB}}$, DCRB remains more difficult to transfer owing to the inclusion of depth-wise convolution, even though at a macroscopic level $\phi(\text{DCRB})\approx\phi(\text{CRB})$.
$\begin{cases}\phi_{1}=\frac{l\cdot e^{k+1}}{f}\cdot\lg\left(\frac{n}{100}\right)\\ \phi_{2}=\frac{l\cdot e^{k+1}\cdot\text{Sigmoid}\left(\frac{n}{100}\right)}{f}=\frac{l\cdot e^{k+1}}{f\cdot\left(1+e^{-n/100}\right)}\\ \phi_{3}=\frac{l\cdot\lg\left(\frac{n}{100}\right)\cdot\text{Swish}(k+1)}{f}=\frac{l\cdot\lg\left(\frac{n}{100}\right)}{f}\frac{k+1}{1+e^{-k-1}}\\ \phi_{4}=\frac{e^{k+1}\cdot\lg(\frac{n}{100})}{f}\cdot l^{2}\\ \phi_{5}=\frac{(l^{2}+1)\cdot(5k^{2}+1)\cdot\sqrt{n+2}}{f}\\ \phi_{6}=\frac{\text{Softplus}(2k+1)\cdot\text{Softplus}(l)\cdot\text{Tanh}(\frac{n}{100})}{f}\end{cases}$ (21)
In summary, an appropriately chosen UAE formulation serves to normalize score ranges, thereby facilitating direct comparisons across modules. Moreover, once Eq.(1) is satisfied, alternative formulations of the UAE do not affect the module universality rankings.
Figure 5: UAE results on plug-and-play/non-plug-and-play modules.
Figure 6: Computational overhead analysis on different plug-and-play modules.
Figure 7: Convergence comparison between plug-and-play modules .
Figure 8: Visual comparison of conventional RB and optimized CRB/DCRB on 148089.png and 97033.png in B100 [36].
Figure 9: Quanlitative comparison of coalmine images for ×$\times 4$ SISR.
Figure 10: Qualitative comparison of remote sensing images for ×$\times 4$ SISR.
4.3 Computational Overhead Analysis
As illustrated in Fig.7, CRB achieves the highest PSNR, while DCRB delivers comparably strong fidelity with the lowest FLOPs, parameter count, and GPU runtime. Furthermore, CRB outperforms RB in PSNR under equivalent parameters (tested on Potsdam [43]). The convergence profiles in Fig.7 indicate that all universal modules exhibit similar training patterns, with CRB attaining marginally faster convergence within the first 20 epochs, a behavior attributable to the accelerated gradient flow afforded by its additional residual branches (Eq.(7)).
4.4 SISR for General Scene Images
Using EDSR [6] as the backbone network, we conduct a combined analysis of module universality and performance, as shown in Table 3. Nearly all of the optimized modules feature better generalizability than baseline RB by slightly sacrificing their universality. Notably, RCAB–despite suffering the greatest universality degradation–fails to achieve the best performance. In contrast, CRB attains the largest generalization gains with minimal impact on universality, marking it as a positively universal module that could be integrated into diverse architectures without additional computational overhead. On the one hand, this finding demonstrates that CRB represents an exceptionally ideal module. On the other hand, this finding indicates that universality and generalizability are not isolated. Although prior work has shown that higher generalization could be attained by sacrificing module transferability—via increased module complexity and representational capacity—their precise interplay remains contingent on the specific architecture of plug-and-play modules.
Figure 11: Hardware deployments and the practical applicability evaluation of optimized modules.
Table 4 summarizes CRB and DCRB performance across multiple SR scales. EDSR augmented by CRB consistently achieves the highest fidelity (PSNR/SSIM), whereas DCRB yields only minor performance degradations—-never exceeding a 5% gap and reaching a minimum of 0.23% under ×$\times 4$ SR on B100. Remarkably, DCRB comprises just 1.73% of the parameter count of CRB and RB (Table 2), yet under certain conditions delivers approximately 99% of the baseline performance.
As illustrated in Fig.8, we present the ×$\times 3$ SR results on the B100 dataset [36] to visually compare CRB (Ours1) and DCRB (Ours2). Both variants achieve marked clarity enhancements over the LR, yielding similar qualitative reconstructions. The residual images-—defined as the per-pixel discrepancy between HR and SR images-—reveal that CRB substantially reduces errors in critical details (e.g., roof eaves in 97033.png). In contrast, DCRB incurs increased residual magnitudes compared to the baseline, particularly within detailed texture areas where errors are more pronounced.
In summary, although DCRB offers superior transferability across model backbones, its positive impact on overall generalizability is limited. In contrast, CRB not only ensures robust plug-and-play universality but also substantially enhances performance across diverse datasets. Accordingly, CRB is best suited for applications with strict performance requirements, whereas DCRB is advantageous for lightweight or resource-constrained deployments.
4.5 SISR For Remote Sensing Images
We employ the Potsdam dataset [43] for both training and testing. Benchmark models include SRCNN [1], SRResNet [18], RCAN [11], EDSR [6], IPG [17], SRFormer [42] as well as their enhanced variants integrating the CRB and DCRB modules.
Table 5 reports the remote-sensing SISR results. Models augmented with CRB consistently outperform their baseline counterparts, demonstrating CRB’s superior universality and its positive impact on model generalizability. In particular, “IPG$+$CRB” achieves the best performance. In contrast, DCRB-—owing to its depthwise separable convolution design-—occasionally underperforms CRB.
Fig.LABEL:remote_sensing_original illustrates the SISR results of remote sensing imagery. Integrating CRB into the model substantially enhances the resolution of ground vehicle, producing more detailed textures and sharper edge contours. Fig.LABEL:remote_sensing_spatial displays the spatial maps of the SISR outputs, facilitating a perceptual comparison of model performance. Both CRB and DCRB effectively suppress bright regions, with the IPG variant augmented by CRB achieving the highest perceptual quality.
4.6 SISR For Industrial Scenes: Applications on Low-Light Mine Imagery Reconstruction
Using the BelT dataset [44], we evaluate the optimized models’ performance in extreme industrial settings (e.g., mine environments), thereby validating the effectiveness of CRB and DCRB.
To reconstruct low-light, low-resolution mine imagery, we cascade KinD [45], the illumination-enhancement module, at the model’s input and perform SISR under normal illumination at the output. Results are summarized in Table 6. Nearly all of the models embedded with CRB/DCRB witness improvements over the baseline while IPG augmented by CRB achieves the best performance.
Fig.LABEL:coalmine_1 provides qualitative evidence: In Image 1, the mechanical structures flanking the belt are reconstructed with sharply defined edges. In Image 2, the transported materials, initially imperceptible, become distinctly visible after reconstruction. In Fig.LABEL:coalmine_2, we plot the pixel intensities along the 125th row of SR outputs. All reconstructed channels show a pronounced increase in amplitude, with CRB yielding marginally higher values than DCRB. This confirms that the optimized modules preserve original image features while amplifying high-frequency details, thus improving both signal fidelity and perceptual realism.
4.7 Hardware Implementations and Extra low-light Applications
In addition to software optimizations, we conduct on-device hardware validation of CRB and DCRB modules using Jetson Nano B01 paired with a CSI camera to assess their practical applicability. The hardware setup is illustrated in Fig.LABEL:hardware.
Fig.LABEL:hardware_results provides hardware-based evidence supporting the effectiveness of proposed modules, where the optimized model successfully reconstructs low-light&low-resolution images that KinD alone cannot reconstruct, thereby markedly enhancing the critical details (e.g., object textures).
From the spectral density plots in Fig.LABEL:hardware_results, it is evident that, unlike the LR image whose spectrum is dominated by an intensely bright central region indicating redundant low-frequency information, the SR image recovers substantially more high-frequency details. This is illustrated by the SR spectral density’s overall brightness extending toward the high-frequency regions and similarity to the HR spectrum. Although a slight haze remains in the reconstructed output, it does not compromise the overall visual quality.
5 Conclusions
In this study, we propose the concept of “Universality” to quantify the transferability of modules. We introduce the UAE to measure module universality and design two optimized variants, CRB and DCRB, on the basis of RB. Analysis on their back propagation process reveals the underlying mechanisms behind these optimization behaviors. Extensive experiments across multiple SISR scenarios (natural scenes, remote sensing data, extreme industrial environments, and real-world applications) demonstrate that our approach outperforms several state-of-the-art models. Moreover, the proposed optimization strategy is applicable to other plug-and-play modules (e.g., RCAB), providing a principled framework for future architectural advances.

### v2

@@
 However, existing researches have predominantly focused on raw performance gains, with little attention paid to quantifying the transferability of architectural components.
-In this paper, we introduce the concept of “Universality” and its associated definitions which extend the traditional notion of “Generalization” to encompass the modules’ ease of transferability, thus revealing the relationships between module universality and model generalizability.
-Then we propose the Universality Assessment Equation (UAE), a metric for quantifying how readily a given module could be transplanted across models.
+In this paper, we introduce the concept of “Universality” and its associated definitions which extend the traditional notion of “Generalization” to encompass the modules’ ease of transferability.
+Then we propose the Universality Assessment Equation (UAE), a metric which quantifies how readily a given module could be transplanted across models and reveals the combined influence of multiple existing metrics on transferability.
 Guided by the UAE results of standard residual blocks and other plug-and-play modules, we further design two optimized modules, Cycle Residual Block (CRB) and Depth-Wise Cycle Residual Block (DCRB).
-Through comprehensive experiments on natural-scene benchmarks, remote-sensing datasets, extreme-industrial imagery and on-device deployments, we demonstrate that networks embedded with the proposed plug-and-play modules outperform several state-of-the-arts, reaching a PSNR enhancement of up to 0.83dB or enabling a 71.3% reduction in parameters with negligible loss in reconstruction fidelity.
-Keywords Generalization; Optimization; Super-resolution; Universality
+Through comprehensive experiments on natural-scene benchmarks, remote-sensing datasets and other low-level tasks, we demonstrate that networks embedded with the proposed plug-and-play modules outperform several state-of-the-arts, reaching a PSNR enhancement of up to 0.83dB or enabling a 71.3% reduction in parameters with negligible loss in reconstruction fidelity.
+Similar optimization approaches could be applied to a broader range of basic modules, offering a new paradigm for the design of plug-and-play modules.
+Keywords Generalization; Optimization; Super-resolution; Universality; Plug-and-Play
 1 Introduction
@@
 To enable a quantitative assessment of module transferability and provide a novel perspective on plug-and-play module optimization, this paper proposes the definition of “Universality” and its associated concepts.
-Subsequently, the Universality Assessment Equation is designed by analyzing the pivotal characteristics of existing plug-and-play modules and other state-of-the-art networks.
-Cycle Residual Blocks (CRB) and Depth-wise Cycle Resiual Blocks (DCRB) are further derived on the basis of RB under the guidance of UAE results.
+Subsequently, the Universality Assessment Equation (UAE) is designed by analyzing the pivotal characteristics of existing plug-and-play modules and other state-of-the-art networks.
+Cycle Residual Blocks (CRB) and Depth-wise Cycle Residual Blocks (DCRB) are further derived on the basis of RB under the guidance of UAE results.
 Experiments demonstrate that models embedded with the optimized modules could achieve a 0.83dB increase in PSNR, or reduce parameters by up to 71.3% with minimal performance loss.
@@
 The inherent optimization mechanisms of CRB and DCRB are also revealed from the standpoint of back propagation process.
-3) The effectiveness of CRB and DCRB is validated through experiments on various SISR tasks ranging from conventional settings to extreme industrial scenarios (e.g., coalmine images), where the optimized modules outperform several state-of-the-arts and significantly enhance the visual quality of reconstructed images.
-Moreover, through on-device deployments, we have validated the practicability of CRB in real-world applications.
+3) The effectiveness of CRB and DCRB is validated through experiments on various SISR tasks, and the UAE optimization strategy is further verified through evaluation on various low-level tasks such as denoising and deblurring.
+Experiments demonstrate that our optimization strategy features strong generalizability across different vision tasks, offering a new paradigm for the optimization of plug-and-play modules.
 2 Related Works
 2.1 Models for Single Image Super-resolution
-Existing SISR networks could be broadly classified into four categories based on their backbone models: CNN-based models ([1, 6, 4]), GNN-based models ([17, 20, 21, 22]), GAN-based models ([23]), and Transformer-based models ([15, 13, 42, 24]).
+Early SISR networks could be broadly classified into four categories based on their backbone models: CNN-based models ([1, 6, 4]), GNN-based models ([22, 25, 26, 27]), GAN-based models ([28]), and Transformer-based models ([20, 13, 46, 29]).
 While highly integrated models could reduce computational costs and improve the performance, they also make it difficult to achieve efficient plug-and-play compatibility across different networks.
 For instance, the RSTB module in SwinIR [13] contains STL layers, which in turn incorporate MSA (Multi-Head Self-Attention) and MLP (Multi-Layer Perceptron).
-The deep integration of these components requires careful structure modifications when transferring the module to other backbones.
+Recent hybrid architectures introduce new paradigms but compromise transferability as well.
+SRMamba-T [14] merges Mamba’s sequential scanning with Transformer’s global attention for long-range dependency modeling.
+However, the fundamental difference between Mamba’s linear scanning mechanism and Transformer’s global self-attention necessitates intricate signal pathways, resulting in considerable architectural coupling.
+HSR-KAN [15] integrates KANs (Kolmogorov-Arnold Networks) with CNN and MLP to reach a trade-off between efficiency and quality , but its specialized fusion structure (e.g., KAN-CAB module) severely limits its task generalizability.
+Although the BUFF (Bayesian Uncertainty Guided Diffusion Model) [16] introduces Bayesian methods into diffusion models to relax the assumption of independent noises, however, the instability of Bayesian training substantially increases model coupling and computational intensity.
+Furthermore, the tight coupling in diffusion models between the forward diffusion and reverse denoising processes precludes any individual component from functioning independently.
+Therefore, while these advances boost performance, their intricate integrations impede module reuse across architectures.
 2.2 Plug-and-Play Modules in SISR Networks
-Existing SISR methods often incorporate various modules into the networks to enhance model performance.
-For instance, RB, introduced in ResNet [18], has been widely adopted in diverse architectures.
-ConvFFN, introduced in ViT [16], is used in the GNN-based IPG [17].
-SRFlow-DA [27] embeds RB directly into the affine injector layer to improve the performance of SRFlow [28].
-These plug-and-play modules significantly enhance model performance but are relatively scarce and difficult to design, since structures of the plug-and-play modules are required to be highly streamlined while preserving an ideal performance.
-Therefore, we attempt to analyze the structure of plug-and-play modules from a universality perspective, thus designing modules with enhanced performance.
+The concept of plug-and-play modules emerges from the recognition that modular components could improve reproducibility across different SISR architectures.
+Early successful examples include the widespread use of RB across diverse networks, from CNN-based EDSR [6] to GAN-based ESRGAN [28], demonstrating the value of structurally simple yet effective components.
+The attention mechanisms represent one of the most successful transferable modules in recent years.
+Beyond RCAB’s [11] channel attention, researchers developed increasingly sophisticated attention variants.
+For instance, MDAB (Multi-scale Dilated Attention Block) [17] offers a lightweight attention mechanism.
+However, its effectiveness critically depends on the pre-processing by a dedicated LRM (Local Residual Module).
+Consequently, transplanting MDAB without its coupled LRM generally leads to performance degradation.
+These plug-and-play modules significantly enhance model performance but remain relatively scarce and challenging to design, since they demand highly streamlined structures while preserving its ideal performance.
+Therefore, analyzing the module structures from a universal perspective is crucial for designing improved plug-and-play modules.
 2.3 Generalization and Universality
-“Generalization” denotes a model’s ability to maintain high performance on test sets distinct from its training data [29].
-IPT [15], for example, utilizes large-scale pre-training to unify three different low-level vision tasks.
-DGUNet [19], based on DNN, designs an interpretable SISR network featuring universality.
-While these studies enhance model performance by improving generalizability, they lack a quantitive evaluation of local module transferability.
-Therefore, this paper introduces the concept of universality to characterize module transferability, thereby facilitating the design of plug-and-play modules that enhance model generalization.
+“Generalization” refers to a model’s capacity to maintain high performance when evaluated on data distinct from its training distribution [34].
+Existing researches exploring “universality” predominantly focus on the developments of unified frameworks and task formulations.
+For instance, IPT [20] leverages large-scale pre-training to unify diverse low-level vision tasks, while DGUNet [24] proposes an interpretable, unified deep neural network architecture.
+Similarly, Uni-COAL[18] introduces a unified framework for Magnetic Resonance Imaging super-resolution.
+Although these approaches enhance model performance by improving framework universality, they have not thoroughly investigated or provided quantitative evaluation for the module transferability.
+Therefore, this paper proposes the concept of universality to characterize the transferability of individual components.
+This quantification facilitates the systematic design of plug-and-play modules aimed at enhancing model generalization and framework universality.
 3 Proposed Methods
-In this section, definitions for “Universality” and its associated concepts are proposed to lay the foundation for a qualitative description of module transferability, namely the Universality Assessment Equation (UAE).
+In this section, definitions for “Universality” and its associated concepts are proposed to lay the foundation for a quantitative description of module transferability, namely the Universality Assessment Equation.
 We further optimize RB into CRB/DCRB through UAE analysis and the optimization mechanism under UAE guidance is elucidated from the perspective of back propagation.
@@
 Concept Subject Object Evaluation Criteria
-Universality Local module Complete model The ease of module transferability
-Generalization Complete model Brand new test data Whether model performance remains ideal
+Universality
+Local
+module
+Complete
+model
+The ease of
+module transferability
+Generalization
+Complete
+model
+Brand new
+test data
+Whether model
+performance remains ideal
 Although universality could describe a module’s intrinsic transferability, it is difficult to characterize its precise impacts on model performance after integration, where directly embedding a plug-and-play module may degrade the performance.
@@
 Differences between the previously proposed plug-and-play modules ([23, 11, 36, 37]) reveal that while more complex blocks could achieve superior generalizability, they also typically incur larger parameter counts and more elaborate computational graphs, all of which may deteriorate the ease of transferability.
-In other words, model’s generalizability could be enhanced by appropriately decreasing the module universality.
+Therefore, generalizability could be enhanced by appropriately decreasing the module universality.
 3.2 Universality Assessment Equation (UAE)
@@
 $\phi=\frac{l\times e^{k+1}\times\lg\left(\frac{n}{100}\right)}{f}$ (2)
-Figure 2: Structures of the proposed CRB and DCRB, where $l$ stands for the UAE variable.
+Figure 2: Overview of the proposed blocks and the UAE optimization strategy.
 3.3 Cycle Residual Block and Depth-wise Separable Cycle Residual Block
@@
 DCRB employs depth-wise separable convolution to reduce the parameters, resulting in a more lightweight and integrated module structure.
-Let $X\in{R}^{B\times C\times H\times W}$ denote the input tensor, where $B$, $C$, $H$ and $W$ correspond to batch size, channel dimensions, height and width, respectively.
+Let $X\in\mathbb{R}^{B\times C\times H\times W}$ denote the input tensor, where $B$, $C$, $H$ and $W$ correspond to batch size, channel dimensions, height and width, respectively.
 The CRB and its depth-wise variant DCRB are then defined as Eq.(3), where Conv($\cdot$) represents standard convolution for CRB and depth-wise separable convolution for DCRB.
@@
 where $L(x)$ stands for losses and $J_{i}$ the Jacobian matrix.
-Let $\mathbf{F}_{l}(\cdot):\mathbb{R}^{d}\xrightarrow{}\mathbb{R}^{d}$ denote the nonlinear mapping (e.g., convolution followed by activation ) in the $l$-th residual block.
+Let $\mathbf{F}_{l}(\cdot)$ denote the nonlinear mapping (e.g., convolution followed by activation ) in the $l$-th residual block.
 Since each block implements $x_{l+1}=x_{l}+\mathbf{F}_{l}(x_{l})$, the total derivative of the loss with respect to initial input $x_{0}$ follows by the repeated application of chain rule:
@@
 $\|(E+A)(E+B)\|_{2}>\|(E+A)\|_{2}.$ (8)
-Figure 3: Experimental validation of Eqs.(5), (7) and (12), where EDSR [6] is chosen as the backbone.
-EDSR$+$CRB attains 31.00 dB PSNR by epoch 200, whereas EDSR$+$RB achieves the same value around epoch 350, confirming CRB’s faster convergence.
-Figure 4: Experimental validation of Eqs.(14) and (16), where EDSR [6] is chosen as the backbone.
-CRBl=16,ε=4 exhibits smooth convergence, whereas CRBl=16,ε=5 fails to converge and possesses a substantially degraded performance, with its loss plateauing at approximately 632.5 during the experiments.
 Let $A,B\in\mathbb{R}^{d\times d}$ satisfy: (1) “$A$ and $B$ are symmetric positive semidefinite matrices”, (2) “$A$ and $B$ commute, i.e.
-$AB=BA$”.
+AB=BAAB=BA”.
 Thus, there exists orthogonal matrix $Q$ satisfying Eq.(9).
@@
 Thus, we have $(E+A)(E+B)$ in this basis equal to $\text{diag}((1+\alpha_{1})(1+\beta_{1}),\cdots,(1+\alpha_{d})(1+\beta_{d}))$. By the definition of $\|\cdot\|_{2}$, we derive Eq.(11).
-[…]
-[…]
+$\|E+A\|_{2}$ $=\underset{i}{\max}(1+\alpha_{i}),\|(E+A)(E+B)\|_{2}$ $=\underset{i}{\max}[(1+\alpha_{i})(1+\beta_{i})].$ (11)
+where $\underset{i}{\max}\beta_{i}\triangleq\beta_{\max}\geq 0$. Finally we arrive at the expression given in Eq.(12).
 $\|(E+A)(E+B)\|_{2}$ $=(1+\alpha_{\max})(1+\beta_{\max})>(1+\alpha_{\max})$ $=\|(E+A)\|_{2}$ (12)
@@
 Treating $F^{\prime}$ as a “small perturbation” aligns with the fact that residual learning ultimately guides deep modules degenerate into identity function.
-In quantum mechanics, Dysen series and Lie-Trotter product formulas [34] use the approximation $e^{\frac{A}{n}}\approx I+\frac{1}{n}A+O(n^{-2})$ retaining merely the linear “$I+$small” at each step, motivating the requirements $\|E+A\|_{2}\geq 1$ in Eq.(12).
-Therefore, even relaxing the symmetric/commuting conditions, Eq.(12) together with Eq.(7) indicate that every added residual branch contributes at least a unit of spectral gain1absent1$\geq$ 1≥ 1 into the back propagation, leading to smoother gradient flow, faster convergence, and better module performance, as is verified through experiments in Fig.3.
+[…]
 Table 2: UAE calculation on different modules, where $\text{UAE}_{i}$ represents six specific forms of UAE and $f=64$. Structure of CRB is chosen as $\varepsilon=2,l=8$.
@@
 $\text{UAE}_{6}(\times 1)$ $\phi_{6}$ 0.08 0.71 0.29 1.20 2.30 0.38 0.16
-Table 3: Quantitative comparison results (PSNR/SSIM) between several universal modules with 4×4$\times$4 × SR scale, trained on DIV2K. Bold texts indicate the best performance.
+Table 3: Quantitative comparison results (PSNR/SSIM) between several universal modules with $4\times$ SR scale, trained on DIV2K.
 Modules RB (baseline) [23] RCAB [11] ConvFFN [21] CRB DCRB
@@
 Urban100 [45] 23.53/0.7597 23.49/0.7414 23.45/0.7502 23.52/0.7615 23.13/0.7312
-[…]
+Table 4: Quantitative comparison (PSNR/SSIM/LPIPS) for RB [6] and CRB/DCRB, using EDSR [6] as the backbone. Oursii represents the incorporation of corresponding module into the baseline.
 SR Scale Model Set5[44] Set14 [43] B100 [40] Urban100 [45]
-×$\times 3$ EDSR (RB) 33.69/0.8844/0.0707 29.87/0.7101/0.0634 28.31/0.8540/0.0980 -/-/-
+$\times 3$ EDSR (RB) 33.69/0.8844/0.0707 29.87/0.7101/0.0634 28.31/0.8540/0.0980 -/-/-
 Ours1(CRB) 33.76/0.8846/0.0684 29.92/0.7114/0.0624 28.33/0.8553/0.0972 -/-/-
 Ours2(DCRB) 32.98/0.8756/0.1248 28.38/0.7072/0.0918 27.56/0.8421/0.1538 -/-/-
-×$\times 4$ EDSR (RB) 31.29/0.8632/0.1092 28.06/0.6961/0.1085 25.95/0.8122/0.2026 23.53/0.7597/0.0247
+$\times 4$ EDSR (RB) 31.29/0.8632/0.1092 28.06/0.6961/0.1085 25.95/0.8122/0.2026 23.53/0.7597/0.0247
 Ours1(CRB) 31.32/0.8648/0.1112 28.49/0.6971/0.1110 25.97/0.8137/0.2008 23.52/0.7615/0.0243
@@
 3.4 Generalized Forms for CRBs and its Designing Principles
-Although Eq.(12) demonstrates that introducing extra residual connections accelerates module convergence without adding extra parameters, we derive an upper bound of $\frac{l}{4}$ on the residual number $\varepsilon$ ($\varepsilon\leqslant\frac{l}{4}$), where $l$ is the CRB’s UAE variable and the constant 4 equals half the number of forward propagation layers per CRB.
-Following Eq.(7), we initially verify that for $\varepsilon=4,l=16$, CRB converges stably. From Fig.2, CRB ($\varepsilon=4,l=16$) could be expressed as Eq.(13)
-$y_{1}=x+\mathbf{F}_{1}(x),y_{2}=y_{1}+\mathbf{F}_{2}(y_{1}),y_{3}=y_{2}+\mathbf{F}_{3}(y_{2}),y_{4}=y_{3}+\mathbf{F}_{4}(y_{3}).$ (13)
-Hence, for a network comprised of $L$ CRBε=4, the back propagation is given by Eq.(14). Fig.LABEL:converge_1 and Fig.LABEL:loss_1 provides the experimental validation for the convergence of Eq.(14).
-$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}\Big{[}(\mathbf{E}+\frac{\partial\mathbf{F}_{4,l}}{\partial y_{3,l}})(\mathbf{E}+\frac{\partial\mathbf{F}_{3,l}}{\partial y_{2,l}})(\mathbf{E}+\frac{\partial\mathbf{F}_{2,l}}{\partial y_{1,l}})(\mathbf{E}+\frac{\partial\mathbf{F}_{1,l}}{\partial x_{l}})\Big{]}=\frac{\partial L}{\partial x_{L}}\prod_{i=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon=4}}\end{aligned}.$ (14)
-However, when an additional residual connection is introduced to Eq.(13), as shown in Eq.(15), the back propagation should be updated as Eq.(16).
-$y_{1}=y_{0}+\mathbf{F}_{1}(y_{0}),y_{2}=y_{1}+\mathbf{F}_{2}(y_{1})+x,y_{3}=y_{2}+\mathbf{F}_{3}(y_{2}),y_{4}=y_{3}+\mathbf{F}_{4}(y_{3}).$ (15)
-$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\Bigg{[}\prod_{l=0}^{L-1}\Big{[}\prod_{i=1}^{4}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})+\Theta_{l}\Big{]}\Bigg{]}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon=5}},\Theta_{l}=\prod_{t=3}^{4}(\mathbf{E}+\frac{\partial\mathbf{F}_{t,l}}{\partial y_{t-1,l}}).$ (16)
-In conjunction with Eq.(12), it is clear that $\|J_{l}^{\text{CRB}_{\varepsilon=5}}\|_{2}>2>\|J_{l}^{\text{CRB}_{\varepsilon=4}}\|_{2}>1$, indicating that $\text{CRB}_{l=16,\varepsilon=5}$ is highly susceptible to trigger gradient exploding.
-Such conclusion derived from theoretical analysis are corroborated by the experimental results presented in Fig.LABEL:converge_2 and Fig.LABEL:loss_2, where $\text{CRB}_{l=16,\varepsilon=5}$ exhibits gradient explosions in first 20 epochs accompanied by a substantial decline in PSNR performance.
-We now generalize the setting to a CRB with $l$ forward propagation layers where $\varepsilon$ residual connections are introduced, as shown in Eq.(17).
-Assuming that the previously stated hypothesis on the upper bound of residual numbers $\varepsilon\leqslant\frac{l}{4}(\varepsilon\in\mathbb{N})$ holds, then the computed results should theoretically indicate that the module converges for $\varepsilon=\frac{l}{4}$ but fails to converge for $\varepsilon=\frac{l}{4}+1$.
-$\begin{cases}y_{1}=x+\mathbf{F}_{1},\cdots,\ y_{\varepsilon}=y_{\varepsilon-1}+\mathbf{F}_{\varepsilon},&\varepsilon=\frac{l}{4}\\ y_{1}=x+\mathbf{F}_{1},\ y_{2}=y_{1}+\mathbf{F}_{2}+x,\cdots&\varepsilon=\frac{l}{4}+1\end{cases}$ (17)
-Referring to Eqs.(14) and (16), we derive the back propagation process ($\varepsilon=\frac{l}{4}$ for Eq.(18) and $\varepsilon=\frac{l}{4}+1$ for Eq.(19)) for the model made up of $L$ generalized CRBs.
-Similar to Eq.(12), it could be easily concluded that $\|J_{l}^{\text{CRB}_{\varepsilon=l/4+1}}\|_{2}>2>\|J_{l}^{\text{CRB}_{\varepsilon=l/4}}\|_{2}>1$.
-Consequently, when $\varepsilon>\frac{l}{4}$, $\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon}}$ exhibits exponential growth with a base larger than 2, leading to gradient explosion and declined module performance.
-$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L}\prod_{i=1}^{\varepsilon}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon}}.$ (18)
-$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\Bigg{[}\prod\limits_{l=0}^{L-1}\Big{[}\prod\limits_{i=1}^{\varepsilon}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})+\Theta_{l}\Big{]}\Bigg{]}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon}},\Theta_{l}=\prod_{t=3}^{\varepsilon}(\mathbf{E}+\frac{\partial\mathbf{F}_{t,l}}{\partial y_{t-1,l}}).$ (19)
-Table 5: Quantitative results on Potsdam remote sensing dataset [43] under ×$\times 4$ SR scale.
+Although Eq.(12) demonstrates that introducing extra residual connections accelerates module convergence without adding extra parameters and is beneficial for performance improvements, we derive an upper bound on the residual number $\varepsilon$ ($\varepsilon\leqslant l/4+\sum\varepsilon_{x}\leqslant l/4+\delta_{\text{safe}},\delta_{\text{safe}}\geq 0$), where $l$ is the UAE variable inferred from computational graphs, $\delta_{\text{safe}}$ is correlated with hyperparameters and experimental settings, and $\varepsilon_{x}$ represents the extra residual connections introduced by inputs.
+Additionally, we theoretically demonstrate that $\sum\varepsilon_{x}$ is inversely proportional to the module’s stability.
+When $\sum\varepsilon_{x}>\delta_{\text{safe}}$, the module undergoes gradient explosion.
+In our ablation study, we reveal the nature of this phenomenon and clarify why the module performance degrades.
+Consider a CRB* as illustrated in Eq.(13), whose $l/4=4,\sum\varepsilon_{x}=|m|+|n|+|z|$. Similar to Eq.(7), we derive the back propagation of a model made up of $n$ CRB*s, as shown in Eq.(14).
+$y_{1}=x+\mathbf{F}_{1},y_{2}=y_{1}+\mathbf{F}_{2}+mx,y_{3}=y_{2}+\mathbf{F}_{3}+nx,y_{4}=y_{3}+\mathbf{F}_{4}+zx$ (13)
+$\frac{\partial L(x)}{\partial x_{0}}$ $=\frac{\partial L(x)}{\partial x_{n}}\prod_{l=0}^{n-1}\Big{[}\prod_{i=1}^{4}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})+m\prod_{i=3}^{4}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})+n(\mathbf{E}+\frac{\partial\mathbf{F}_{4,l}}{\partial y_{3,l}})+z\mathbf{E}\Big{]}$ (14)
+$=\frac{\partial L(x)}{\partial x_{n}}\prod_{l=0}^{n-1}J_{l}^{\text{CRB*}}$
+It is obvious that $1+\sum\varepsilon_{x}<\|J_{l}^{\text{CRB*}}\|_{2}<2+\sum\varepsilon_{x}$, where $\sum\varepsilon_{x}=|m|+|n|+|z|$.
+To reveal the threshold conditions that prevent modules from experiencing gradient explosion, we consider an extreme condition where $\sum\varepsilon_{x}=\delta_{\text{safe}}$, leading to: $\|\partial L(x)/\partial x_{0}\|_{2}\in\Big{(}(1+\delta_{\text{safe}})^{n},(2+\delta_{\text{safe}})^{n}\Big{)}$.
+Therefore, to minimize the risk of gradient explosion, it is necessary to set $\delta_{\text{safe}}=0$.
+When $\delta_{\text{safe}}>0$, various factors (e.g., hyperparameters) collectively govern the module’s convergence, and the risk of gradient explosion escalates as $\delta_{\text{safe}}$ increases.
+To generalize our conclusion, we further consider a CRB with $l$ forward propagation layers where $\varepsilon$ residual connections are introduced, as shown in Eq.(15).
+$y_{1}=x+\mathbf{F}_{1},y_{2}=y_{1}+\mathbf{F}_{2}+c_{1}x,y_{3}=y_{2}+\mathbf{F}_{3}+c_{2}x,\cdots,\ y_{\varepsilon}=y_{\varepsilon-1}+\mathbf{F}_{\varepsilon}+c_{\varepsilon-1}x$ (15)
+Referring to Eqs.(13) and (14), we derive the back propagation process for the model made up of $L$ generalized CRBs.
+By defining $D_{k}=\partial y_{k}/\partial x$, the recurrence relation for $D_{k}$ is established as Eq.(16).
+$D_{k}=(\mathbf{E}+\frac{\partial\mathbf{F}_{k}}{\partial y_{k-1}})D_{k-1}+c_{k-1}\mathbf{E},k\geq 2$ (16)
+By calculating Eq.(16) , the general form of $D_{k}$ goes as:
+$D_{k}=\prod_{i=1}^{k}(\mathbf{E}+\frac{\partial\mathbf{F}_{i}}{\partial y_{i-1}})+\sum_{j=1}^{k-1}c_{j}\prod_{i=j+2}^{k}(\mathbf{E}+\frac{\partial\mathbf{F}_{i}}{\partial y_{i-1}})+c_{k-1}\mathbf{E}$ (17)
+When setting $k=l$, we derive the Jacobian matrix of generalized CRB:
+$J_{l}=D_{k=l}=\prod_{i=1}^{l}(\mathbf{E}+\frac{\partial\mathbf{F}_{i}}{\partial y_{i-1}})+\sum_{j=1}^{l-1}c_{j}\prod_{i=j+2}^{l}(\mathbf{E}+\frac{\partial\mathbf{F}_{i}}{\partial y_{i-1}})+c_{l-1}\mathbf{E}$ (18)
+Similarly, $1+\sum_{j=1}^{l-1}|c_{j}|=1+\sum\varepsilon_{x}<\|J\|_{2}<2+\sum\varepsilon_{x}$.
+Consequently, to most straightforwardly ensure that CRB could enhance performance with the least probability of encountering gradient explosion, $\delta_{\text{safe}}$ can be simply set to zero.
+Table 5: Quantitative results on Potsdam remote sensing dataset [47] under $\times 4$ SR scale.
 Hyper-parameter settings of IPG and SRFormer are scaled down to accommodate limited computational resources and accelerate training, while EDSR and SRResnet–both featuring moderate parameter counts–retain their full hyper-parameter configurations.
@@
 IPG (with DCRB) 32.85 0.9029 0.0565
-Table 6: Quantitative results on BelT coalmine dataset [44] under ×$\times 4$ SR scale. Bold texts indicate the best performance.
-Models PSNR(dB) SSIM LPIPS
-EDSR (CVPR 2017) [6] 28.06 0.7551 0.1154
-SRFormer (ICCV 2023) [42] 29.76 0.7906 0.1017
-IPG (CVPR 2024) [17] 29.95 0.7921 0.1028
-EDSR (with CRB) 28.54 0.7604 0.1122
-IPG (with CRB) 30.03 0.7930 0.1010
-IPG (with DCRB) 29.97 0.7928 0.0999
-SRFormer (with CRB) 29.81 0.7884 0.1042
-SRFormer (with DCRB) 29.80 0.7913 0.1033
-In summary, the generalized CRB converges stably and achieves optimal performance only when $\varepsilon\leqslant\frac{l}{4}$ is satisfied, where $\varepsilon$ is the number of residual connections and $l$ the number of UAE propagation layers (Eq.(1)).
 4 Experimental analysis
@@
 Although RCAB [11] possesses larger $n$ and $l$ compared to RSTB [13], the elevated $k$ in RSTB deteriorates its transferability.
-This finding aligns with empirical evidence: RSTB integrates three nested submodules—MSA, MLP, and STL—whereas RCAB employs only a CALayer.
+This finding aligns with empirical evidence: RSTB integrates three nested submodules: MSA, MLP, and STL, whereas RCAB employs only the channel attention.
 Hence, effective migration of RSTB requires the concurrent adaptation of multiple components.
@@
 Figure 6: Visual comparison of conventional RB and optimized CRB/DCRB on 148089.png and 97033.png in B100 [40].
-Figure 9: Quanlitative comparison of coalmine images for ×$\times 4$ SISR.
-Figure 10: Qualitative comparison of remote sensing images for ×$\times 4$ SISR.
+Figure 7: Qualitative comparison of remote sensing images for $\times 4$ SISR.
+Figure 8: Visualization of the reconstructed results.
+“Model+” denotes insertion of the original plug-and-play module, “Model++” the UAE-optimized module, and “Model+++” the module further refined by both UAE and CRB.
+All embedded modules are in accordance with Table 7.
+4.3 Comparison Between UAE and Other Metrics
+Metrics Evaluation Object Measurement Focus Unit
+Parameters Model, Module Storage requirements, Model capacity K($10^{3}$), M($10^{6}$)
+FLOPs Model, Module Computational overhead, Inference speed GFLOPs
+PSNR, SSIM Model output Reconstruction quality, Structural Similarity dB, Dimensionless ([0,1])
+Parameter efficiency Model, Module Performance per parameter Metric per parameter
+GPU time Model, Module Inference time on hardware ms
+UAE (ours) Module Structural flexibility, Transferability Dimensionless ($0,+\infty]$)
+Table 6: Comparison between UAE and other existing metrics.
+Figure 9: Spearman Correlation Analysis on UAE.
+“GPU” represents modules’ inference time on GPU, “PE” stands for parameter efficiency.
+Fig.LABEL:Spearman_alpha_only, LABEL:Spearman_beta_only, LABEL:Spearman_theta_only present the Spearman correlation analysis between each single component of UAE and other metrics, aiming to investigate which aspects of existing metrics are captured by $k,n,l$ in UAE.
+As shown in Table 6, UAE differs from conventional metrics in several aspects.
+While existing metrics such as PSNR and GPU time primarily focus on performance evaluation at the model/module level, UAE is specifically designed to evaluate the structural flexibility and transferability of individual modules.
+This enables UAE to capture an essential yet previously unmeasured aspect of module design: its ability to migrate across different networks.
+Additionally, given that module transferability is affected by multiple factors, Fig.9 employs the Spearman Correlation Analysis to demonstrate that UAE is a composite metric integrating several metrics.
+Each single factor $\alpha,\beta,\theta$ exhibits moderate to strong correlation with certain existing metrics.
+For instance, $\alpha(l)$ correlates highly with CPU time and FLOPs, while $\beta(k)$ has a moderate correlation with parameter efficiency.
+$\beta$ exhibits a strong negative correlation with PSNR, because modules with very deep nesting levels fail to function independently and must operate within a complete model.
+However, UAE evaluates the performance of an isolated module to investigate its transferability.
+As a result, for modules with large $k$, the PSNR performance tends to decrease as the nesting depths increases, leading to a pronounced negative correlation.
+The full UAE metric (Fig.LABEL:Spearman_UAE) results demonstrate that UAE is not dominated by a single factor (e.g., 0.72 with parameters, 0.5 with GPU time), but rather offers a more balanced assessment that captures aspects not reflected by existing metrics alone.
+Therefore, UAE is a novel metric that reveals the combined influence of multiple existing metrics on transferability, an effect not captured by any individual metric.
 4.4 Computational Overhead Analysis
-As illustrated in Fig.7, CRB achieves the highest PSNR, while DCRB delivers comparably strong fidelity with the lowest FLOPs, parameter count, and GPU runtime.
+As illustrated in Fig.5, CRB achieves the highest PSNR, while DCRB delivers comparably strong fidelity with the lowest FLOPs, parameter count, and GPU time.
 Furthermore, CRB outperforms RB in PSNR under equivalent parameters (tested on Potsdam [47]).
-The convergence profiles in Fig.7 indicate that all universal modules exhibit similar training patterns, with CRB attaining marginally faster convergence within the first 20 epochs, a behavior attributable to the accelerated gradient flow afforded by its additional residual branches (Eq.(7)).
+The convergence profiles in Fig.5 indicate that all universal modules exhibit similar training patterns, with CRB attaining marginally faster convergence within the first 20 epochs, a behavior attributable to the accelerated gradient flow afforded by its additional residual branches.
 4.5 SISR for General Scene Images
@@
 In contrast, CRB attains the largest generalization gains with minimal impact on universality, marking it as a positively universal module that could be integrated into diverse architectures without additional computational overhead.
-On the one hand, this finding demonstrates that CRB represents an exceptionally ideal module.
-On the other hand, this finding indicates that universality and generalizability are not isolated.
-Although prior work has shown that higher generalization could be attained by sacrificing module transferability—via increased module complexity and representational capacity—their precise interplay remains contingent on the specific architecture of plug-and-play modules.
-Figure 11: Hardware deployments and the practical applicability evaluation of optimized modules.
+This finding demonstrates that CRB represents an exceptionally ideal module and indicates that universality and generalizability are not isolated.
+Although prior work has shown that higher generalization could be attained by sacrificing module transferability via an increased complexity and representational capacity, their precise interplay remains contingent on the specific architecture of plug-and-play modules.
 Table 4 summarizes CRB and DCRB performance across multiple SR scales.
-EDSR augmented by CRB consistently achieves the highest fidelity (PSNR/SSIM), whereas DCRB yields only minor performance degradations—-never exceeding a 5% gap and reaching a minimum of 0.23% under ×$\times 4$ SR on B100.
+EDSR augmented by CRB consistently achieves the highest fidelity (PSNR/SSIM), whereas DCRB yields only minor performance degradations—-never exceeding a 5% gap and reaching a minimum of 0.23% under $\times 4$ SR on B100.
 Remarkably, DCRB comprises just 1.73% of the parameter count of CRB and RB (Table 2), yet under certain conditions delivers approximately 99% of the baseline performance.
-As illustrated in Fig.8, we present the ×$\times 3$ SR results on the B100 dataset [36] to visually compare CRB (Ours1) and DCRB (Ours2).
+As illustrated in Fig.6, we present the $\times 3$ SR results on the B100 test set to visually compare CRB (Ours1) and DCRB (Ours2).
 Both variants achieve marked clarity enhancements over the LR, yielding similar qualitative reconstructions.
-The residual images-—defined as the per-pixel discrepancy between HR and SR images-—reveal that CRB substantially reduces errors in critical details (e.g., roof eaves in 97033.png).
+The residual images reveal that CRB substantially reduces errors in critical details (e.g., roof eaves in 97033.png).
 In contrast, DCRB incurs increased residual magnitudes compared to the baseline, particularly within detailed texture areas where errors are more pronounced.
@@
 Models augmented with CRB consistently outperform their baseline counterparts, demonstrating CRB’s superior universality and its positive impact on model generalizability.
-In particular, “IPG$+$CRB” achieves the best performance.
-In contrast, DCRB-—owing to its depthwise separable convolution design-—occasionally underperforms CRB.
+In particular, “IPG++CRB” achieves the best performance.
+In contrast, owing to DCRB’s depthwise separable convolution design, DCRB occasionally underperforms CRB.
 Fig.LABEL:remote_sensing_original illustrates the SISR results of remote sensing imagery.
@@
 Both CRB and DCRB effectively suppress bright regions, with the IPG variant augmented by CRB achieving the highest perceptual quality.
-4.6 SISR For Industrial Scenes: Applications on Low-Light Mine Imagery Reconstruction
-Using the BelT dataset [44], we evaluate the optimized models’ performance in extreme industrial settings (e.g., mine environments), thereby validating the effectiveness of CRB and DCRB.
-To reconstruct low-light, low-resolution mine imagery, we cascade KinD [45], the illumination-enhancement module, at the model’s input and perform SISR under normal illumination at the output.
-Results are summarized in Table 6.
-Nearly all of the models embedded with CRB/DCRB witness improvements over the baseline while IPG augmented by CRB achieves the best performance.
-Fig.LABEL:coalmine_1 provides qualitative evidence: In Image 1, the mechanical structures flanking the belt are reconstructed with sharply defined edges.
-In Image 2, the transported materials, initially imperceptible, become distinctly visible after reconstruction.
-In Fig.LABEL:coalmine_2, we plot the pixel intensities along the 125th row of SR outputs.
-All reconstructed channels show a pronounced increase in amplitude, with CRB yielding marginally higher values than DCRB.
-This confirms that the optimized modules preserve original image features while amplifying high-frequency details, thus improving both signal fidelity and perceptual realism.
-4.7 Hardware Implementations and Extra low-light Applications
-In addition to software optimizations, we conduct on-device hardware validation of CRB and DCRB modules using Jetson Nano B01 paired with a CSI camera to assess their practical applicability.
-The hardware setup is illustrated in Fig.LABEL:hardware.
-Fig.LABEL:hardware_results provides hardware-based evidence supporting the effectiveness of proposed modules, where the optimized model successfully reconstructs low-light&low-resolution images that KinD alone cannot reconstruct, thereby markedly enhancing the critical details (e.g., object textures).
-From the spectral density plots in Fig.LABEL:hardware_results, it is evident that, unlike the LR image whose spectrum is dominated by an intensely bright central region indicating redundant low-frequency information, the SR image recovers substantially more high-frequency details.
-This is illustrated by the SR spectral density’s overall brightness extending toward the high-frequency regions and similarity to the HR spectrum.
-Although a slight haze remains in the reconstructed output, it does not compromise the overall visual quality.
+4.7 UAE Optimization on Other Basic Modules and Low-Level Tasks
+Deblurring Models GOPRO
+PSNR SSIM LPIPS
+FMANet (CVPR 2024) 27.57 0.8339 0.0783
+PPA (ICME 2024) + FMANet 27.71 0.8383 0.0725
+$\text{PPA}^{\text{*}}$ + FMANet 27.69 0.8398 0.0719
+$\text{PPA}^{\text{**}}$ + FMANet 27.72 0.8387 0.0699
+Deraining Models Rain200H
+PSNR SSIM LPIPS
+NeRD-Rain (CVPR 2024) 30.13 0.9145 0.0430
+ELAB (ECCV 2022) + NeRD-Rain 30.19 0.9140 0.0428
+$\text{ELAB}^{\text{*}}$ + NeRD-Rain 30.23 0.9143 0.0426
+$\text{ELAB}^{\text{**}}$ + NeRD-Rain 30.42 0.9161 0.0422
+Denoising Models SIDD-Medium
+PSNR SSIM LPIPS
+AFM (CVPR 2024) 29.63 0.5843 0.0093
+Agent-Attention (ECCV 2024) + AFM 30.61 0.6180 0.0091
+$\text{Agent-Attention}^{\text{*}}$ + AFM 31.76 0.6920 0.0089
+$\text{Agent-Attention}^{\text{**}}$ + AFM 32.56 0.7490 0.0077
+Joint Denoising & Low-Light-Enhancement Models VILNC
+PSNR SSIM LPIPS
+$\text{ZeroIG}^{{\dagger}}$ (CVPR 2024) 10.17 0.2952 0.4708
+ARConv (CVPR 2025) + $\text{ZeroIG}^{{\dagger}}$ 12.13 0.3353 0.3809
+$\text{ARConv}^{\text{*}}$ + $\text{ZeroIG}^{{\dagger}}$ 12.23 0.4346 0.3932
+$\text{ARConv}^{\text{**}}$ + $\text{ZeroIG}^{{\dagger}}$ 12.50 0.3806 0.3911
+Table 7: UAE optimization on different modules and performance evaluation of the optimized models on various low-level tasks.
+$\text{Module}^{\text{*}}$” represents the module optimized by UAE, and $\text{Module}^{\text{**}}$” represents the CRBs embedded on the basis of UAE optimization.
+Among the optimized variants, the structural parameters of ZeroIG are moderately reduced to adapt to limited computing resources (marked as $\text{ZeroIG}^{{\dagger}}$), resulting in reduced model performance.
+However, model performance could still be compared by measuring the relative values of metrics between models.
+Firstly, we briefly summarize the models and corresponding UAE optimization methods for the four tasks.
+Since our UAE optimization stays only within the module, our methods do not affect the inherent plug-and-play characteristics of the modules.
+We only select appropriate places to embed our optimized components into the framework without changing models’ core architectures.
+FMANet (Flow-Guided Dynamic Filtering and Iterative Feature Refinement with Multi-Attention Network) [51] comprises a multi-layer convolutional feature extractor followed by point-wise mappings, emphasizing per-pixel enhancement via stacked convolution blocks.
+PPA (Parallelized Patch-aware Attention) [52] is a multi-branch convolutional block fused with local/global attention and spatial attention, used for multi-scale context aggregation.
+During our UAE optimization process, we augment PPA with channel expansion, parallel depth-wise branches and branch fusion followed by spatial attention to increase the representational capacity.
+CRB is further inserted after $\text{PPA}^{\text{*}}$ output as a post-refinement stage.
+NeRD-Rain (Neural Representations for Image Deraining) [53] is a multi-level architecture built from window processing, using local filters and stacked convolutions for deraining.
+ELAB (Efficient Long-rang Attention Block) [54] is a block combining local filtering and gated multi-window self-attention (GMSA), excelling at local detail refinement.
+When designing $\text{ELAB}^{\text{*}}$, we add a lightweight local depth-wise branch and SE-style channel reweighting to better use input features, while the GMSA structure is fully preserved.
+Further, CRBs are applied after $\text{ELAB}^{\text{*}}$ as a bidirectional residual refinement for detail correction.
+AFM (Adversarial Frequency Mixup Framework) [55] is a lightweight DnCNN-style backbone, consisting of multiple convolution layers that perform per-pixel noise restoration.
+Agent-Attention [56] derives adaptive pooled agent tokens and injects agent information into features.
+We optimize Agent-Attention by increasing heads and agent numbers, and we further embed CRBs inside the agent wrapper, where the attention output is projected and then passed through CRB before normalization and residual additions.
+ZeroIG (Zero-Shot Illumination Guided Framework) [57] is a decomposition-based architecture with Enhancer and Denoising modules arranged in a cascaded fashion, which could perform denoising and low-light-enhancement simultaneously.
+ARConv (Adaptive Rectangular Convolution) [58] is an adaptive resampling convolution module which fixes the drawbacks of convolution operations within a confined square window.
+We further wrap the ARConv with channel expansion and multiple parallel branches to better leverage the input features.
+CRBs are embedded in ZeroIG’s high-frequency branch (H3).
+To be specific, after the denoising step, CRBs are applied for detail refinement.
+Table 7 demonstrates that models embedded with the plug-and-play modules exhibit better performance than the baselines, and after our UAE optimization, the performance has generally been further improved.
+Additionally, models embedded with the modules that are optimized by both UAE and CRB have achieved new SOTAs on nearly all the datasets.
+From a qualitative perspective, as shown in Fig.
+8, images restored by the optimized models also have better visual quality than the baselines, and most of the details in LQ (Low Quality) images have been recovered.
+Furthermore, as shown in the deraining part of Fig.
+8, although the performance of NeRD-Rain+ is slightly degraded after directly embedding the plug-and-play module into NeRD-Rain, after UAE optimization and CRB enhancement, however, the model performance still surpasses the baseline, providing direct evidence for the effectiveness of our UAE optimization strategy.
+Therefore, our optimization strategy is not only applicable to SR tasks, but also other low-level tasks such as denoising, deblurring, and deraining.
+4.8 Ablation Studies
+4.8.1 Sensitivity Quantification of UAE Parameters
+$\alpha(l)$ $\beta(k)$ $\theta(n)$ RB RCAB ConvFFN RSTB GAL DCRB CRB
+0.063 0.234 0.094 0.172 0.328 0.125 0.125
+0.011 0.028 0.028 0.061 0.061 0.028 0.030
+0.045 0.050 0.035 0.046 0.043 0.017 0.045
+0.033 0.087 0.062 0.180 0.169 0.030 0.033
+0.179 0.743 0.211 0.505 0.902 0.138 0.359
+0.046 0.413 0.165 0.675 1.289 0.220 0.091
+0.131 1.309 0.372 1.984 3.543 0.244 0.262
+Table 8: Variable Ablation of UAE, where $\phi$3 is specified as the UAE form and the module parameters refer to Table 2.
+A logarithmic derivative-based approach and elasticity coefficients are proposed to quantify the variable sensitivity.
+For a general UAE form where $\phi=\Pi_{i}f_{i}(x_{i})$, the logarithmic differentiation goes as Eq.(21)
+$d\ln{\phi}=\sum_{i}\frac{\partial\ln{f_{i}}}{\partial\ln{x_{i}}}d\ln{x_{i}}$ (21)
+The sensitivity of $x_{i}$ is defined as the absolute elasticity and its normalized form, as shown in Eq.(22).
+$S_{i}$ represents the percentage change in $\phi$ caused by a 1% change in $x_{i}$.
+$x_{i}$ and $f_{i}$ respectively stand for the variables and functions in UAE.
+For instance, $\{<x_{\alpha},f_{\alpha}>\}=\{<l,\alpha(l)>\}$.
+$S_{i}=\left|\frac{\partial\ln{\phi_{i}}}{\partial\ln{x_{i}}}\right|=\left|\frac{x_{i}}{f_{i}}\cdot\frac{\partial f_{i}}{\partial x_{i}}\right|,\hat{S_{i}}=\frac{S_{i}}{S_{\alpha}+S_{\beta}+S_{\theta}},i\in\{\alpha,\beta,\theta\}$ (22)
+Additionally, we employ the Sharply value to measure the contribution of each variable, which differs from the sensitivity.
+Sharply value is a concept from cooperative game theory that allocates the total contribution of each participant by averaging their marginal contributions across all combinations.
+Given that UAE features three variables, the Sharply equations are defined as Eq.(23).
+$\begin{cases}C_{\alpha}=\frac{v(\alpha)-v(\varnothing)+v(\alpha,\beta)-v(\beta)+v(\alpha,\theta)-v(\theta)}{3}\\ C_{\beta}=\frac{v(\beta)-v(\varnothing)+v(\alpha,\beta)-v(\alpha)+v(\beta,\theta)-v(\theta)}{3}\\ C_{\theta}=\frac{v(\theta)-v(\varnothing)+v(\theta,\beta)-v(\beta)+v(\alpha,\theta)-v(\alpha)}{3}\end{cases}$ (23)
+Figure 10: Variable Ablation of UAE.
+Fig.LABEL:Variable_Ablation_1 demonstrates the effects of $x,<x,y>,<x,y,z>(x,y,z\in\{k,n,l\})$ on UAE calculation, and Fig.LABEL:Variable_Ablation_2 depicts the variable sensitivity via elasticity coefficients.
+Fig.LABEL:Variable_Abaltion_3 describes the absolute variable contributions on UAE.
+[…]
+4.8.2 Nature of CRB Gradient Explosion and Experimental Validation on the $\mathbf{\varepsilon}$ Constraints
+$l$ $\varepsilon$ $\varepsilon\leqslant(l/4+\delta_{\text{safe}})$ $\mathcal{L}_{1}^{\text{min}}$ Set5 (PSNR/SSIM) Set14 (PSNR/SSIM) B100 (PSNR/SSIM)
+8 1 7.31 27.39/0.8610 24.17/0.7470 23.96/0.7129
+2 7.47 27.36/0.8607 24.15/0.7469 23.95/0.7128
+3 15.39 21.46/0.5703 20.58/0.5203 21.93/0.5220
+12 2 7.57 27.53/0.8643 24.25/0.7503 23.94/0.7126
+3 7.50 27.32/0.8601 24.13/0.7468 23.98/0.7143
+4 27.06 17.50/0.3236 17.46/0.3062 19.18/0.3427
+16 3 8.00 27.41/0.8622 24.18/0.7486 23.95/0.7133
+4 7.77 27.46/0.8629 24.21/0.7492 23.96/0.7135
+5 101.92 10.53/0.0320 11.22/0.0349 13.00/0.0629
+Table 9: Ablation study of $\varepsilon$.
+Models are trained for 400 epochs on DIV2K dataset, EDSR is used as the backbone.
+$\delta_{\text{safe}}$ is set to zero during the experiments.
+$\mathcal{L}_{1}^{\text{min}}$ represents the minimum L1 loss during training.
+As shown in Table 9, when $\varepsilon$ exceeds the boundary $l/4+\delta_{\text{safe}}$, module performance significantly deteriorates, and when $\varepsilon\leqslant l/4+\delta_{\text{safe}}$, module performance remains remarkably similar.
+Therefore, excessive residual connections do not further enhance model performance.
+Figure 11: Gradient norm heatmaps of CRB variants during the training process. $l$ represents the module depth and $\varepsilon$ the number of residual connections.
+Fig.14 reveals a critical phenomenon validating the constraint $l/4+\delta_{\text{safe}}$.
+For a cascade of $m=5$ CRBs, when $\varepsilon\leqslant l/4+\delta_{\text{safe}}$ is satisfied, the gradient norms $\|\nabla_{w}L\|$ for the convolution layers within the initial CRB (closest to the input) are smaller than those within the terminal CRB (closest to the output).
+This is because $J_{l}^{\text{CRB}_{\varepsilon}}$ for each block satisfies $\|J_{l}^{\text{CRB}_{\varepsilon}}\|_{2}>1$ but still remains bounded, preventing gradient vanishing and explosion.
+The total gradients at the input of block $k$ can be expressed as:
+$\|\nabla_{x_{k}}L\|_{2}\propto\|\nabla_{x_{m}}L\|_{2}\cdot\prod_{j=k}^{m-1}\|J_{j}^{\text{CRB}_{\varepsilon}}\|_{2}$ (24)
+While the residual structure $\|J_{j}^{\text{CRB}_{\varepsilon}}\|_{2}>1$ mitigates vanishing gradient, the cumulative product over blocks still induces controlled attenuation towards the input, resulting in $\|\nabla_{w_{1}}L\|<\|\nabla_{w_{m}}L\|$, as shown in Fig.LABEL:CRB_8l_2epsilon, LABEL:CRB_12l_3epsilon, LABEL:CRB_16l_4epsilon.
+This reflects stable, well-behaved back propagation.
+When $\varepsilon>l/4+\delta_{\text{safe}}$, however, the gradient norms $\|\nabla_{w}L\|$ for convolution layers within the initial CRB becomes larger than those within the terminal CRB.
+This is because violating $\varepsilon\leqslant l/4+\delta_{\text{safe}}$ triggers $\|J_{l}^{\text{CRB}_{\varepsilon}}\|>2$.
+Therefore, gradient at the input of block $k$ scales exponentially as:
+$\|\nabla_{x_{k}}L\|_{2}\propto\|\nabla_{x_{m}}L\|_{2}\cdot\prod_{j=k}^{m-1}\|J_{j}^{\text{CRB}_{\varepsilon}}\|_{2}>\|\nabla_{x_{m}}L\|_{2}\cdot 2^{(m-k)}$ (25)
+where the exponential amplification dominates the gradient flow.
+Given that the input block undergoes amplification over Jacobian multiplications, it is obvious that $\|\nabla_{w_{1}}L\|_{2}>\|\nabla_{w_{m}}L\|_{2}$, manifesting the gradient explosion predicted theoretically.
+Fig.LABEL:explosion1, LABEL:explosion2, LABEL:explosion3 reveal the explosion’s origin in the earliest layers in CRB1, and this abnormal gradient behaviour directly explains the significant performance degradation observed in Table 9.
 5 Conclusions
@@
 Analysis on their back propagation process reveals the underlying mechanisms behind these optimization behaviors.
-Extensive experiments across multiple SISR scenarios (natural scenes, remote sensing data, extreme industrial environments, and real-world applications) demonstrate that our approach outperforms several state-of-the-art models.
-Moreover, the proposed optimization strategy is applicable to other plug-and-play modules (e.g., RCAB), providing a principled framework for future architectural advances.
+Extensive experiments across multiple SISR scenarios and other low-level tasks demonstrate that our approach outperforms several state-of-the-art models.
+Moreover, the proposed optimization strategy is applicable to other basic modules, providing a principled framework for future architectural advances.
+Author Contributions
+Haotong Cheng: Conceptualization, formal analysis, methodology, software, writing-original draft.
+Zhiqi Zhang: Investigation, validation, visualization.
+Hao Li and Xinshang Zhang: Resources, supervision, writing-review & editing.
+Conflicts of Interest
+The authors declare no conflicts of interest.
+Data Availability Statement
+This study uses publicly available datasets.
+The datasets for training are available via the following links: DIV2K(https://data.vision.ee.ethz.ch/cvl/DIV2K/), BelT(https://github.com/CUMT-AIPR-Lab/CUMT-AIPR-Lab).
+Datasets for evaluation are accessible in the following link: Set5, Set14, B100, Urban100(https://gitcode.com/Resource-Bundle-Collection/efef9/).
+Other Low-Level datasets could be found directly in their corresponding references.
+Experimental datas are provided within the article.